# Orchestrated Robust Controller for the Precision Control of Heavy-duty Hydraulic Manipulators

### v2

@@ VI Stability Analysis
In the previous section, the modeling and control of the HHM were established. Fig. 6 demonstrates the scheme of the proposed approach with details of subsystem-based modeling and control. Fig. 6 shows that the required force designed in the rigid body subsystem is transmitted to the actuator subsystem to be generated. The connections between rigid body subsystems and actuator subsystems are established by means of VPFs. In this section, the stability of the entire system under the proposed controller is analyzed. The adaptation functions for handling parameter uncertainties in rigid body parts and actuator parts are derived to achieve SGUUB.
@@
^˙$^{}^{A}\dot{\hat{\mathcal{L}}}=\frac{1}{\gamma}\,^{A}\hat{\mathcal{L}}\,\left({}^{A}\mathcal{S}-\gamma_{0}\,^{A}\hat{\mathcal{L}}\right)\,^{A}\hat{\mathcal{L}}$ (105)
@@
is virtually stable in the sense of Definition 4, where $A\in S$ and $S=\left\{P_{1},P_{p2},B_{0j},B_{1j},B_{3j},B_{4j},G_{i}\right\}$ for $i=1,2,3$ and $j=1,2$. Moreover, ${}^{A}\mathcal{S}$ is a unique symmetric matrix. Let the non-negative accompanying function for the rigid body part be chosen in the sense of Definition 3 as ${\nu}_{1}$. Then, the time derivative of ${\nu}_{1}$ is:
@@
VII Results and Discussion
In this section, the simulation and experimental results are provided to evaluate the performance of the proposed controller. First, the simulation results of comparing the presented ORC method with three different controllers, including the original VDC, proportional-differential (PD) controller, and the state-of-the-art controller, adaptive dynamic surface controller with funnel control and neural networks (ADSC-FC-NN) proposed in [17]. The simulation analysis is performed for the base joint of the HHM, shown in Fig. 1, which is subjected to deadzone in valves and backlash in the rack and pinion mechanism. Then, the performance of the proposed method is implemented to control the real-world commercial heavy-duty hydraulic manipulator.
@@ VII-A Performance comparison in Simulation
VII-B Experimental Set-up
Fig. 8 demonstrates the experimental set-up for performance evaluation of the controller. In the following, the hardware components and properties are provided:
@@ VII-B Experimental Set-up
In order to implement the designed controller, only encoder and pressure sensor data are utilized. Using the pressure sensor to compute the piston force in (86) for automation of heavy-duty manipulators is a more practical solution since retrofitting the pressure sensors is much easier than retrofitting the piston load cell. In order to perform numerical differentiation for $\frac{d}{dt}({{}^{A}}V_{r})$ in (65) and $\dot{f}_{pr}$ and $\dot{x}_{p}$ in (100), the finite difference method introduced in [39] is used with a low-pass filter. The control goal is to achieve five given set points in Cartesian space with the largest magnitude of $0.8\,m$, $1\,m$, and $1\,m$ in $x$, $y$, and $z$ directions, respectively, in different orientations. The Euler angles with the $XYZ$ convention are used to represent the orientation of the end-effector with respect to the base, and the orientation error is computed based on the rotation matrix. Then, a smooth fifth-order trajectory generator [40] is utilized to produce a smooth trajectory between the set points for a given execution time, $t_{f}$. The (46) and (47) are used to compute the joint’s desired values. Fig. 9 demonstrates the motion of the robot.
@@
Fig. 8: Experimental set-up
VII-C Experimental Result
In this section, the experimental results of implementing the proposed controller on an industrial 6-DoF HHM are provided. The result of the proposed method is compared to that of the original VDC controller to evaluate the control performance more accurately. Table II provides the control gains utilized in experiments. For the RBFNNs, the Gaussian activation function is defined as $\Psi(\chi)=exp([-(\chi-c_{\mathbf{j}})^{T}(\chi-c_{\mathbf{j}})/(b_{\mathbf{j}}^{2})])$, with $c_{\mathbf{j}}$ and $b_{\mathbf{j}}$ denoting the center and width of the neural cell in $\mathbf{j}$th unit, with $\mathbf{j}=\bar{n}_{A}$ for rigid body subsystem, and $\mathbf{j}=\bar{n}_{a}$ for actuator subsystem. In this study the value for $c_{\mathbf{j}}$ is randomly selected in $[-1,1]\times[-1,1]\times[-1,1]$ and $[-1,1]\times[-1,1]\times[-1,1]\times[-1,1]$ for $\chi_{A}$ and $\chi_{a}$, respectively, with $b_{\bar{n}_{a}}=0.5$ and $b_{\bar{n}_{A}}=5$ with the aggregation of 270 nodes. Additionally, prior to inputting the vectors into RBFNN, we normalize all input features to the range of $[-1,1]$. This step ensures that all input dimensions are treated uniformly and that the RBFNNs can effectively cover the input space without the risk of incompatibility.
In Fig. 10, the performance of the proposed method and the original VDC in tracking a desired trajectory is compared. As demonstrated, both controllers performed excellently in following the desired trajectory. However, from Fig. 11, it can be concluded that the proposed controller achieved a much better result by handling the uncertainties in the model and considering the input nonlinearities. As mentioned previously, the base and wrist actuators (objects 1 and 3) suffer from compound input nonlinearities. Such uncertainty along with the model uncertainties lead to the steady-state error of 0.13, -0.6, -0.7, and 0.07 degree in $\zeta_{1}$, $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, respectively, using VDC controller. By employing the presented controller, the errors are reduced to -0.04, -0.01, -0.02, and -0.01 degrees, respectively. In addition to the steady state errors, the proposed method achieved a considerably lower transient error, showing the perfect performance of the designed method in tackling unknown uncertainties and compound input nonlinearities. Fig. 12, on the other hand, depicts the angular velocity of each joint along with the desired values, displaying that the ORC ended up with much smoother tracking without jumps in velocities.
@@ VII-C Experimental Result
Finally, experimental results are juxtaposed to those reported in [5], utilizing $\rho$ value as a metric for comparative assessment. Table IV exhibits the computed performance index value for all the works along with its normalization based on the end-effector DoF (EE-DoF), indicating number of DoFs controlled in Cartesian space with 3 DoF for position and 3-DoF for orientation. As can be seen, the proposed method has achieved a better $\rho$ value in comparison to other works with full pose control, demonstrating its excellent behavior in Cartesian space motion control.
Considering all the experimental results provided in this section, along with the comparison to the original VDC approach, the validity of the theoretical results is verified. As shown, tackling unknown model and actuator uncertainties with RBFNNs and handling compound input nonlinearities with an adaptive deadzone-backlash controller significantly improved the tracking performance, both in transient and steady-state. As it was mathematically proved in Theorem 3 and experimentally shown in this section, the designed controller ensured SGUUB stability for the system and converged the tracking error to a substantially small neighborhood of origin.
@@
VIII Conclusion
@@
In this study, an orchestrated robust controller is designed to achieve a precise and robust performance. For this purpose, VDC was employed as the baseline controller. VDC decomposes the entire complex system into subsystems of the rigid body and actuators by means of VCP. Then, the controller design and stability analysis are accomplished at the subsystem level and expanded to the entire system by means of VPFs and virtual stability. A novel way of incorporating RBFNNs and adaptive inverse deadzone-backlash controllers into each subsystem resulted in the orchestrated robust controller approach. Considering all unknown model and actuator uncertainties, compound input nonlinearities, and unknown disturbances, the SGUUB stability is achieved for the first time in the context of VDC. All theoretical results are validated via extensive simulation and experimental results concerning the highly nonlinear, six DoF HHM. The simulation results demonstrated that the proposed controller outperformed the original VDC, PD, and the state-of-the-art method in the presence of compound input nonlinearities. Moreover, the experimental results demonstrate that the proposed method tackles unknown model and actuator uncertainties and reduces the impact of compound input nonlinearities in real-world applications, resulting in considerably lower transient and steady-state tracking errors and lower actuator force errors in comparison to the original VDC for a given power. Additionally, the proposed method ended up with a precise tracking performance in Cartesian space in comparison to the original VDC. All the analyses, including sensitivity to control gain, tracking performance in joint space and Cartesian space, comparison to the state-of-the-art controller, performance evaluation in the sense of $\rho$ value, and implementation on the real-world generic manipulator, display the universality of the proposed method in theory and applications.
@@
$\begin{split}{\nu}_{1}&=\sum_{A\in S}\dfrac{1}{2}\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)^{T}\,{\rm M_{\rm A}}\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)\\ &+\,\sum_{A\in S}\gamma\mathcal{D}_{F}(\mathcal{L}_{A}\rVert\hat{\mathcal{L}}_{A})+\frac{1}{2}tr({{}^{A}\tilde{W}^{T}}\,{{}^{A}\Gamma^{-1}}{{}^{A}\tilde{W}})\\ &+\frac{1}{2\,{{}^{A}}\pi}{{}^{A}\tilde{\varepsilon}^{T}}\,{{}^{A}\tilde{\varepsilon}}.\end{split}$ (117)
@@
$\begin{split}{\nu}_{1}&\leq\mu_{1}(\sum_{A\in S}\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)^{T}\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)\\ &+tr({{}^{A}\tilde{W}}^{T}{{}^{A}\tilde{W}})+{{}^{A}\tilde{\varepsilon}^{T}}{{}^{A}\tilde{\varepsilon}})\end{split}$ (118)
with $\mu_{1}=\max\left\{\gamma,\{\lambda_{max}(M),0.5\left(\lambda_{max}(\Gamma)\right),0.5{{}^{A}}\pi^{-1}\}_{A\in S}\right\}$, where $\lambda_{max}(.)$ is the maximum eigenvalue of the matrix. Taking the time derivative of (117) with using (116), defining ${}^{A}s=\,^{A}Y^{T}\,(^{A}{V}_{r}-\,^{A}{V})$, and ${}^{A}\tilde{\phi}^{T}\,{{}^{A}s}=tr(\,{{}^{A}\tilde{\mathcal{L}}}\,{{}^{A}\mathcal{S}})$, exploiting $(^{A}{V}_{r}-\,^{A}{V})^{T}\,{{}^{A}}\tilde{W}^{T}\Psi(\chi_{A})=tr((^{A}{V}_{r}-\,^{A}{V})\,\Psi^{T}(\chi_{A})\,{{}^{A}}\tilde{W})$, and replacing from (105)-(107), one can achieve:
$\begin{split}\dot{{\nu}}_{1}&=\,\sum_{A\in S}\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)^{T}\,({{}^{A}F_{r}^{*}}-{{}^{A}F^{*}})\\ &-\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)^{T}K_{A}\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)\\ &-\gamma_{0}tr(\hat{\mathcal{L}}_{A}\,\tilde{\mathcal{L}}_{A})-{{}^{A}}\tau_{0}tr(\,{{}^{A}\hat{W}^{T}}{{}^{A}\tilde{W}})-{{}^{A}}\pi_{0}{{{}^{A}}\tilde{\varepsilon}^{T}}{{}^{A}}\hat{\varepsilon}.\end{split}$ (119)
The first term in (119) is the VPF of the corresponding frame $\left\{A\right\}$. In order to expand the VPFs, we need to use the Definition 2 for each frame in $\left\{S\right\}$. Starting from $\left\{P_{1}\right\}$ and recalling (3), (44), and (71), we have:
@@
In the same sense, we can write for all other frames in $\left\{S\right\}$:
@@
where $VPFs=p_{P_{1}}+p_{{P_{p2}}}+p_{G_{1}}+p_{G_{2}}+p_{G_{3}}-p_{E_{1}}-p_{E_{2}}-p_{E_{3}}-p_{E_{4}}+\sum_{j=1,2}(f_{cjr}-f_{cj})(\dot{x}_{pjr}-\dot{x}_{pj})$. Additionally, $\mu=\mu_{2}/\mu_{1}$, where $\mu_{2}=\min\left\{\gamma_{0},\{\lambda_{min}(^{A}K),0.5{{}^{A}}\tau_{0},0.5{{}^{A}}\pi_{0}^{-1}\}_{A\in S}\right\}$, and $\mu_{0}=\sum_{A\in S}0.5\gamma_{0}tr(\mathcal{L}_{A}\,\mathcal{L}_{A})+0.5{{}^{A}}\tau_{0}tr(\,{{{}^{A}}W^{T}}{{}^{A}W})+0.5{{}^{A}}\pi_{0}{{{}^{A}}\varepsilon^{T}}{{}^{A}}\varepsilon$.

### v4

@@ VI Stability Analysis
In the previous section, the modeling and control of the HHM were established. Fig. 6 demonstrates the scheme of the proposed approach with details of subsystem-based modeling and control. It is shown that the required force designed in the rigid body subsystem is transmitted to the actuator subsystem to be generated. The connections between rigid body subsystems and actuator subsystems are established by means of VPFs. In this section, the stability of the entire system under the proposed controller is analyzed. The adaptation functions for handling parameter uncertainties in rigid body parts and actuator parts are derived to achieve SGUUB.
@@
^˙$^{}^{A}\dot{\hat{\mathcal{L}}}=\frac{1}{\gamma}\,^{A}\hat{\mathcal{L}}\,\left({}^{A}\mathcal{I}-\gamma_{0}\,^{A}\hat{\mathcal{L}}\right)\,^{A}\hat{\mathcal{L}}$ (105)
@@
is virtually stable in the sense of Definition 4, where $A\in\Upsilon$ and $\Upsilon=\left\{P_{1},P_{p2},B_{0j},B_{1j},B_{3j},B_{4j},G_{i}\right\}$ for $i=1,2,3$ and $j=1,2$. Moreover, ${}^{A}\mathcal{I}$ is a unique symmetric matrix, defined Appendix -D. Let the non-negative accompanying function for the rigid body part be chosen in the sense of Definition 3 as ${\nu}_{1}$. Then, the time derivative of ${\nu}_{1}$ is:
@@
VII Simulation Results
In this section, the simulation results are provided to evaluate the performance of the proposed controller in comparison to other methods. Three different controllers are employed as follow; original VDC, proportional-differential (PD) controller, and the state-of-the-art controller, adaptive dynamic surface controller with funnel control and neural networks (ADSC-FC-NN) proposed in [20]. The simulation analysis is performed for the base joint of the HHM, shown in Fig. 1, which is subjected to deadzone in valves and backlash in the rack and pinion mechanism.
@@ VII Simulation Results
VIII Experimental Results
In this section, the performance of the proposed controller is evaluated in real-world scenarios. Fig. 8 demonstrates the experimental set-up for performance evaluation of the controller. In the following, the hardware components and properties are provided:
@@ VIII Experimental Results
In order to implement the designed controller, only encoder and pressure sensor data are utilized. Using the pressure sensor to compute the piston force in (86) for automation of heavy-duty manipulators is a more practical solution since retrofitting the pressure sensors is much easier than retrofitting the piston load cell. In order to perform numerical differentiation for $\frac{d}{dt}({{}^{A}}V_{r})$ in (65) and $\dot{f}_{pr}$ and $\dot{x}_{p}$ in (100), the finite difference method introduced in [44] is used with a low-pass filter. A smooth fifth-order trajectory generator [45] is utilized to produce a smooth trajectory between the set points for a given execution time, $t_{f}$. The (46) and (47) are used to compute the joint’s desired values.
To ensure safe and effective gain tuning for the 6-DoF manipulator, we employed a manual approach guided by safety precautions. Initially, the manipulator was moved to a specific position using joystick commands, and the controller was activated with low gains to verify feedforward term functionality and system stability. Gains were then incrementally increased to improve positional accuracy while maintaining stability. Finally, during trajectory execution, the gains were fine-tuned in real-time to achieve satisfactory performance without compromising the safety of the system.
Fig. 8: Experimental set-up
@@
VIII-A Motion Performance Analysis
In this section, the experimental results of implementing the proposed controller on an industrial 6-DoF HHM are provided. The control goal is to achieve five given set points in Cartesian space with the largest magnitude of $0.8\,m$, $1\,m$, and $1\,m$ in $x$, $y$, and $z$ directions, respectively, in different orientations. The Euler angles with the $XYZ$ convention are used to represent the orientation of the end-effector with respect to the base, and the orientation error is computed based on the rotation matrix. Fig. 9 demonstrates the motion of the robot.
The result of the proposed method is compared to that of the original VDC controller to evaluate the control performance more accurately. Table II provides the control gains utilized in experiments. For the RBFNNs, the Gaussian activation function is defined as $\Psi(\chi)=exp([-(\chi-c_{\mathbf{j}})^{T}(\chi-c_{\mathbf{j}})/(b_{\mathbf{j}}^{2})])$, with $c_{\mathbf{j}}$ and $b_{\mathbf{j}}$ denoting the center and width of the neural cell in $\mathbf{j}$th unit, with $\mathbf{j}=\bar{n}_{A}$ for rigid body subsystem, and $\mathbf{j}=\bar{n}_{a}$ for actuator subsystem. In this study the value for $c_{\mathbf{j}}$ is randomly selected in $[-1,1]\times[-1,1]\times[-1,1]$ and $[-1,1]\times[-1,1]\times[-1,1]\times[-1,1]$ for $\chi_{A}$ and $\chi_{a}$, respectively, with $b_{\bar{n}_{a}}=0.5$ and $b_{\bar{n}_{A}}=5$ with the aggregation of 270 nodes. Additionally, prior to inputting the vectors into RBFNN, we normalize all input features to the range of $[-1,1]$. This step ensures that all input dimensions are treated uniformly and that the RBFNNs can effectively cover the input space without the risk of incompatibility.
In Fig. 10, the performance of the proposed method and the original VDC in tracking a desired trajectory is compared. As demonstrated, both controllers performed excellently in following the desired trajectory. However, from Fig. 11, it can be concluded that the proposed controller achieved a much better result by handling the uncertainties in the model and considering the input nonlinearities. As mentioned previously, the base and wrist actuators (objects 1 and 3) suffer from compound input nonlinearities. Such nonliniearities along with the model uncertainties lead to the steady-state error of 0.13, -0.6, -0.7, and 0.07 degree in $\zeta_{1}$, $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, respectively, using VDC controller. By employing the presented controller, the errors are reduced to -0.04, -0.01, -0.02, and -0.01 degrees, respectively. In addition to the steady state errors, the proposed method achieved a considerably lower transient error, showing the perfect performance of the designed method in tackling unknown uncertainties and compound input nonlinearities. Fig. 12, on the other hand, depicts the angular velocity of each joint along with the desired values, displaying that the ORC ended up with much smoother tracking without jumps in velocities.
@@ VIII-A Motion Performance Analysis
So far, the provided experimental results have thoroughly evaluated the position tracking accuracy of the proposed controller in both joint and Cartesian spaces. The following key insights demonstrate the exceptional performance and impact of the method:
  • The steady-state errors in both the base and wrist joints, specifically in $\zeta_{1}$, $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, which were previously suffering from backlash nonlinearity, have shown substantial improvement. The errors have been reduced by an outstanding 69.23%, 98.33%, 97.14%, and 85.71%, respectively. This demonstrates the effectiveness of the controller in overcoming complex nonlinearities.
  • For each individual joint (from the base to the wrist), a significant improvement in maximum tracking error has been achieved, with reductions of 42.5%, 74.19%, 36.84%, 49.71%, 69.23%, and 20.76%. These reductions are a testament to the robustness of the controller and its precision despite of unknown uncertainties and model complexities.
  • These remarkable improvements in joint space directly translate into a 70% reduction in the total end-effector position tracking error, and a 69.1% reduction in the total end-effector orientation error, as shown in Fig. 15. This is a dramatic improvement in the overall system accuracy, enabling precision operations in complex environments.
  • The sensitivity of the controller to different gains are evaluated, revealing that with 70%, 90%, and 100% of the gains applied, the total RMSE of all joints are 0.455 degrees, 0.27 degrees, and 0.157 degrees, respectively. These results highlight the controller’s robustness, demonstrating that while performance slightly decreases with changes in gain, the system maintains its overall reliability and accuracy.
  • All these performance improvements have been achieved while maintaining the same level of control effort, highlighting the efficiency and practicality of the proposed controller. This ensures that the system can operate at a high performance without requiring excessive computational resources or additional power consumption (almost 70% of CPU were used to ensure real-time computation).
The performance index defined in (1) is a valid and insightful indicator of the controller’s performance. Unlike conventional metrics, this index incorporates both the velocity of the system and the tracking error, offering a more comprehensive perspective on system behavior. Achieving low tracking errors at lower velocities is relatively straightforward; however, increased system velocity excites higher frequencies of the unmodeled dynamics, significantly challenging the controller’s ability to maintain precise motion tracking. The $\rho$ index is uniquely suited to evaluate the performance under these conditions, as it effectively captures the velocity-dependent accuracy of the system. A critical aspect of $\rho$ is its ability to explain the velocity at which the tracking error is achieved. Lower $\rho$ values consistently represent superior tracking accuracy at a given velocity. However, to ensure fairness and relevance in comparisons, the DoFs of the system must be accounted for. An increase in system DoFs inherently imposes greater demands on the controller, and thus, direct comparisons of $\rho$ values across methods without normalization can lead to misleading conclusions. It was also reported in[40] that the established $\rho$ with two DoF was smaller than with three DoF, indicating the effect of nonlinearity in control performance.
@@
To address this, this study normalizes the performance index relative to the controlled DoFs in the Cartesian space for the end-effector (EE-DoF). For example, the maximum EE-DoF is six, encompassing three DoFs for position and three for orientation. Table IV presents the computed $\rho$ values from this study alongside those reported in [8], including their normalization based on EE-DoF.
The results in Table IV clearly highlight the superiority of the proposed method. Even when compared to approaches with fewer EE-DoFs, which inherently benefit from less complexity, the proposed method demonstrates a better normalized performance index. This underscores the importance of normalization in fairly assessing controller performance. Remarkably, the proposed method achieves precise motion tracking for a 6-DoF HHM, despite facing significant challenges such as unknown model uncertainties, input nonlinearities, and system complexities.
Considering all the experimental results provided in this section, along with the comparison to the original VDC approach and other existing methods, the validity of the theoretical results is verified. As shown, tackling unknown model and actuator uncertainties with RBFNNs and handling compound input nonlinearities with an adaptive deadzone-backlash controller significantly improved the tracking performance, both in transient and steady-state. As it was mathematically proved in Theorem 3 and experimentally shown in this section, the designed controller ensured SGUUB stability for the system and converged the tracking error to a substantially small neighborhood of origin.
Fig. 16: Performance analysis in the presence of external disturbances and uncertainties. Each image shows the transition from one set point to another each of which is one meter.
VIII-B Disturbance Rejection Performance Analysis
In this section, we evaluate the robustness of the proposed controller against unknown external disturbances and uncertainties. To achieve this, we attach a two-meter-long rope to the end-effector of the HHM while it operates in full-pose control mode, as shown in Fig. 16. Along this rope, we secure two 50 kg weights, each positioned one meter apart. The experiment begins with the end-effector following a predefined trajectory in the y-direction. At the midpoint of this motion, the weights are suddenly added, creating an unexpected disturbance that the controller must compensate for in real time. The sequence of steps are demonstrated in Fig. 16a-e. Initially, this additional mass acts as an unknown disturbance, introducing oscillations due to its swinging motion during the lift. This dynamic interference negatively impacts the controller’s performance, temporarily altering the system dynamics and requiring rapid compensation. However, as the motion continues, the disturbance becomes an unknown uncertainty, requiring the controller to adapt and maintain precise tracking performance. This setup effectively simulates real-world scenarios where the system encounters unforeseen dynamic changes. The experiments are conducted at three different end-effector velocities: 0.6 m/s, 0.7 m/s, and 0.8 m/s along the direction of motion. This variation enables a more comprehensive evaluation of the proposed controller’s ability to handle external disturbances across practical working velocities.
As shown in Fig. 17a, the proposed controller successfully maintains precise path tracking in the y-direction despite external disturbances, while also regulating motion in the x- and z-directions with high accuracy. Additionally, Figs. 17b and c present the position and orientation tracking errors, respectively. The y-direction error, which is most affected by external disturbances, remains within 2 cm, with minor peaks caused by disturbance interference. Despite this challenge, the controller demonstrates strong robustness while maintaining excellent tracking performance. Moreover, the errors in the x- and z-directions are minimal, highlighting the controller’s ability to effectively isolate disturbances and maintain stability across all motion axes. The total RMSE of position and orientation error were 3.24 mm and 0.15 degree for this experiment. Fig. 17d showcases the velocity tracking performance, where the proposed ORC method precisely follows the desired velocity trajectory with RMSE of 0.02 cm/s, ensuring smooth motion of the heavy manipulator. Finally, Fig. 17e presents the voltage command applied to the manipulator’s valves, revealing a smooth control signal. This further underscores the practicality and real-world applicability of the proposed controller, proving its effectiveness in handling dynamic disturbances while achieving high-performance motion control.
Fig. 17: Disturbance-rejection performance of the proposed ORC controller with an end-effector velocity of approximately 0.6 m/s. Dashed lines are desired values, while solid lines are actual values.
Fig. 18: Disturbance-rejection performance of the proposed ORC controller with an end-effector velocity of approximately 0.7 m/s. Dashed lines are desired values, while solid lines are actual values.
Figs. 18 and 19 present the experimental results for end-effector velocities of 0.7 m/s and 0.8 m/s, respectively. These results provide a striking demonstration of the controller’s ability to handle extreme dynamic conditions. As expected, increasing the end-effector velocity amplifies the effect of external disturbances. However, the proposed controller not only stabilized the system despite the highly nonlinear and unpredictable behavior induced by the swinging loads but also achieved low tracking errors, with peak deviations contained around just 2 cm. The overall RMSE for position and orientation tracking at 0.7 m/s were 3.6 mm and 0.15 degrees, respectively, while at 0.8 m/s, they remained at 5.12 mm and 0.17 degrees. These results highlight the unprecedented accuracy and robustness of the controller, especially considering the extreme challenges posed by such high-velocity motion in the direction of gravity for a heavy manipulator with almost 5 meters workspace. The controller’s ability to maintain such superior tracking accuracy under these conditions is a testament to its performance, making it an feasible solution for precision motion control in real-world, high-impact applications.
Fig. 19: Disturbance-rejection performance of the proposed ORC controller with an end-effector velocity of approximately 0.8 m/s. Dashed lines are desired values, while solid lines are actual values.
VIII-C Discussion
The proposed control method has been successfully implemented on a real-world 6-DoF HHM using a Beckhoff industrial PC with a 1ms sampling time, achieving efficient real-time performance with approximately 70% CPU utilization. While the hybrid nature of the controller, combining model-based (VDC) and non-model-based (RBFNNs) schemes, ensures high precision, increasing the system’s DoF could lead to higher computational demands and potentially limit real-time implementability on less capable hardware. Additionally, higher DoF systems would require more controller gains, making the tuning process more arduous. However, the controller’s robustness to gain variations, which primarily impacts tracking accuracy rather than stability, helps mitigate this challenge. These considerations highlight common challenges for high-precision control systems, yet the demonstrated performance of the proposed method validates its practical applicability.
IX Conclusion
@@
In this study, an orchestrated robust controller is designed to achieve a precise and robust performance. For this purpose, VDC was employed as the baseline controller. VDC decomposes the entire complex system into subsystems of the rigid body and actuators by means of VCP. Then, the controller design and stability analysis are accomplished at the subsystem level and expanded to the entire system by means of VPFs and virtual stability. A novel way of incorporating RBFNNs and adaptive inverse deadzone-backlash controllers into each subsystem resulted in the orchestrated robust controller approach. Considering all unknown model and actuator uncertainties, compound input nonlinearities, and unknown disturbances, the SGUUB stability is achieved for the first time in the context of VDC. All theoretical results are validated via extensive simulation and experimental results concerning the highly nonlinear, six DoF HHM. The simulation results demonstrated that the proposed controller outperformed the original VDC, PD, and the state-of-the-art method in the presence of compound input nonlinearities. Moreover, the experimental results demonstrate that the proposed method tackles unknown model and actuator uncertainties and reduces the impact of compound input nonlinearities in real-world applications, resulting in considerably lower transient and steady-state tracking errors and lower actuator force errors in comparison to the original VDC for a given power. Additionally, the proposed method ended up with a precise tracking performance in Cartesian space in comparison to the original VDC. The robustness of the scheme, moreover, is examined in presence of unknown external disturbances and its excellent performance is demonstrated. All the analyses, including sensitivity to control gain, tracking performance in joint space and Cartesian space, comparison to the state-of-the-art controller, performance evaluation in the sense of $\rho$ value, disturbance rejection performance, and implementation on the real-world generic manipulator, display the universality of the proposed method in theory and applications.
Potential future works could focus on extending the proposed control algorithm to tasks involving both free motion and interaction with the environment, such as pick-and-place operations. This could include the design of an impedance control framework to improve performance in scenarios requiring environmental contact. Additionally, exploring techniques to enhance energy efficiency would make the approach more suitable for mobile platforms. Another promising direction would be the integration of intelligent task planning algorithms, such as reinforcement learning, to enable high-level decision-making and broaden the applicability of the proposed control method in industrial settings.
@@
$\begin{split}{\nu}_{1}&=\sum_{A\in\Upsilon}\dfrac{1}{2}\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)^{T}\,{\rm M_{\rm A}}\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)\\ &+\,\sum_{A\in\Upsilon}\gamma\mathcal{D}_{F}(\mathcal{L}_{A}\rVert\hat{\mathcal{L}}_{A})+\frac{1}{2}tr({{}^{A}\tilde{W}^{T}}\,{{}^{A}\Gamma^{-1}}{{}^{A}\tilde{W}})\\ &+\frac{1}{2\,{{}^{A}}\pi}{{}^{A}\tilde{\varepsilon}^{T}}\,{{}^{A}\tilde{\varepsilon}}.\end{split}$ (117)
@@
$\begin{split}{\nu}_{1}&\leq\mu_{1}(\sum_{A\in\Upsilon}\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)^{T}\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)\\ &+tr({{}^{A}\tilde{W}}^{T}{{}^{A}\tilde{W}})+{{}^{A}\tilde{\varepsilon}^{T}}{{}^{A}\tilde{\varepsilon}}+tr(\tilde{\mathcal{L}}_{A}\,\tilde{\mathcal{L}}_{A}))\end{split}$ (118)
with $\mu_{1}=\max\left\{\gamma,\{\lambda_{max}(M),0.5\left(\lambda_{max}(\Gamma)\right),0.5{{}^{A}}\pi^{-1}\}_{A\in\Upsilon}\right\}$, where $\lambda_{max}(.)$ is the maximum eigenvalue of the matrix. Taking the time derivative of (117) with using (116), defining ${}^{A}\mathbf{s}=\,^{A}Y^{T}\,(^{A}{V}_{r}-\,^{A}{V})$, and ${}^{A}\tilde{\phi}^{T}\,{{}^{A}\mathbf{s}}=tr(\,{{}^{A}\tilde{\mathcal{L}}}\,{{}^{A}\mathcal{I}(\mathbf{s})})$ (${}^{A}\mathcal{I}(\mathbf{s})$ defined in Appendix -D), exploiting $(^{A}{V}_{r}-\,^{A}{V})^{T}\,{{}^{A}}\tilde{W}^{T}\Psi(\chi_{A})=tr((^{A}{V}_{r}-\,^{A}{V})\,\Psi^{T}(\chi_{A})\,{{}^{A}}\tilde{W})$, and replacing from (105)-(107), one can achieve:
$\begin{split}\dot{{\nu}}_{1}&=\,\sum_{A\in\Upsilon}\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)^{T}\,({{}^{A}F_{r}^{*}}-{{}^{A}F^{*}})\\ &-\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)^{T}K_{A}\,\left({{}^{A}{V}_{r}}-{{}^{A}{V}}\right)\\ &-\gamma_{0}tr(\hat{\mathcal{L}}_{A}\,\tilde{\mathcal{L}}_{A})-{{}^{A}}\tau_{0}tr(\,{{}^{A}\hat{W}^{T}}{{}^{A}\tilde{W}})-{{}^{A}}\pi_{0}{{{}^{A}}\tilde{\varepsilon}^{T}}{{}^{A}}\hat{\varepsilon}.\end{split}$ (119)
The first term in (119) is the VPF of the corresponding frame $\left\{A\right\}$. In order to expand the VPFs, we need to use the Definition 2 for each frame in $\left\{\Upsilon\right\}$. Starting from $\left\{P_{1}\right\}$ and recalling (3), (44), and (71), we have:
@@
In the same sense, we can write for all other frames in $\left\{\Upsilon\right\}$:
@@
where $VPFs=p_{P_{1}}+p_{{P_{p2}}}+p_{G_{1}}+p_{G_{2}}+p_{G_{3}}-p_{E_{1}}-p_{E_{2}}-p_{E_{3}}-p_{E_{4}}+\sum_{j=1,2}(f_{cjr}-f_{cj})(\dot{x}_{pjr}-\dot{x}_{pj})$. Additionally, $\mu=\mu_{2}/\mu_{1}$, where $\mu_{2}=\min\left\{\gamma_{0},\{\lambda_{min}(^{A}K),0.5{{}^{A}}\tau_{0},0.5{{}^{A}}\pi_{0}^{-1}\}_{A\in\Upsilon}\right\}$, and $\mu_{0}=\sum_{A\in\Upsilon}0.5\gamma_{0}tr(\mathcal{L}_{A}\,\mathcal{L}_{A})+0.5{{}^{A}}\tau_{0}tr(\,{{{}^{A}}W^{T}}{{}^{A}W})+0.5{{}^{A}}\pi_{0}{{{}^{A}}\varepsilon^{T}}{{}^{A}}\varepsilon$.